OpenStudy (anonymous):

An online store that has been successfully growing on its initial angel investment and revenue wants to invest $5 million to expand the business. The bank is willing to lend the business this money at a 10 percent interest rate over an eight-year term. Calculate the monthly payment and explain what the business must be able to do with this money in order for this to be a smart business decision.

OpenStudy (anonymous):

At simple interest: 10% = 5,000, 000 x 10/100 = 5,000,000 x (1/10) =$500,000 spread over 8 years = 8 x $62,500 monthly interest payment = $62,500/12=$5208.33 Smart decision: has to at least break-even, if not generate more than the total interest repayment.